# Why In-Context Learning Transformers are Tabular Data Classifiers

### v1

@@ A.6 TabZilla Further Results
Models Rank N. Accuracy
min max mean median mean median
TabForestPFN - Fine-tune 1 24 8.0 6.0 0.680 0.663
TabPFN (retrained) - Fine-tune 1 26 8.3 7.8 0.678 0.700
CatBoost 1 22 8.9 7.8 0.676 0.663
XGBoost 1 23 9.2 8.2 0.674 0.671
TabPFN (original) - Fine-tune 1 25 9.3 8.0 0.670 0.677
TabForestPFN - Zero-shot 1 26 10.6 10.0 0.633 0.650
TabForest - Fine-tune 1 24 10.6 9.5 0.650 0.667
TabPFN (original) - Zero-shot 1 25 10.9 10.2 0.632 0.606
LightGBM 1 26 11.1 11.0 0.646 0.653
RandomForest 1 25 11.3 11.0 0.636 0.637
Resnet 1 26 11.9 10.0 0.602 0.613
NODE 1 26 12.1 12.0 0.611 0.596
SAINT 1 26 12.2 12.8 0.600 0.614
SVM 1 25 12.5 13.2 0.589 0.573
FT-Transformer 1 23 12.6 12.5 0.599 0.601
DANet 3 25 14.5 15.0 0.596 0.608
MLP-rtdl 1 26 15.8 17.8 0.514 0.527
TabForest - Zero-shot 3 25 15.9 17.5 0.571 0.583
TabPFN (retrained) - Zero-shot 3 25 15.9 17.5 0.571 0.583
STG 1 26 16.1 17.2 0.495 0.472
LinearRegression 1 26 17.5 19.5 0.459 0.431
MLP 1 26 17.7 20.0 0.475 0.417
TabNet 2 26 18.0 19.0 0.503 0.490
DecisionTree 1 26 18.7 21.0 0.413 0.372
KNN 2 26 19.4 21.5 0.399 0.372
VIME 1 26 21.8 23.8 0.295 0.221
In the TabZilla main results Table 3, we have shown the five highest performing methods. The full results are shown in Table 6, which includes all ICL-transformer variants and methods implemented by the TabZilla authors. Because the rank is calculated over all included methods, which ICL-transformer variants we include might change the results. Therefore, we check if the results are the same if we use calculate the rankings one ICL-transformer at the time.
@@
TabForestPFN - Fine-tune 1 18 5.8 4.2 0.680 0.663
CatBoost 1 15 6.2 5.2 0.676 0.663
XGBoost 1 16 6.3 5.0 0.674 0.671
LightGBM 1 19 7.6 6.0 0.646 0.653
RandomForest 1 18 7.9 8.0 0.636 0.637
NODE 1 19 8.4 8.0 0.611 0.596
Resnet 1 19 8.4 8.0 0.602 0.613
SAINT 1 19 8.5 7.2 0.600 0.614
SVM 1 18 8.6 8.0 0.589 0.573
FT-Transformer 1 16 8.8 8.5 0.599 0.601
DANet 2 19 10.0 10.0 0.596 0.608
MLP-rtdl 1 19 11.4 12.0 0.514 0.527
STG 1 19 11.5 12.0 0.495 0.472
LinearRegression 1 19 12.3 13.8 0.459 0.431
MLP 1 19 12.6 14.0 0.475 0.417
TabNet 1 19 12.8 13.8 0.503 0.490
DecisionTree 1 19 13.3 14.0 0.413 0.372
KNN 2 19 13.9 15.0 0.399 0.372
VIME 1 19 15.7 17.0 0.295 0.221
@@
TabPFN (original) - Zero-shot 1 19 7.4 6.0 0.632 0.606
TabPFN (original) - Fine-tune 1 19 6.4 5.0 0.670 0.677
TabPFN (retrained) - Zero-shot 2 19 10.7 11.8 0.571 0.583
TabPFN (retrained) - Fine-tune 1 19 5.9 5.0 0.678 0.700
TabForest - Zero-shot 2 19 10.7 11.8 0.571 0.583
TabForest - Fine-tune 1 19 7.4 7.0 0.650 0.667
TabForestPFN - Zero-shot 1 19 7.2 6.0 0.633 0.650
TabForestPFN - Fine-tune 1 18 5.8 4.2 0.680 0.663
@@ A.7 WhyTrees Further Results
The main results in Figure 4 report the normalized accuracy aggregated over all datasets. In Figure 8 we show the comparison between ICL-transformers on all 23 datasets. In Figure 9 we show the colored results on each individual dataset where we compare fine-tuned TabForestPFN with all the benchmark methods.
Figure 8: Comparison of TabForest, TabPFN and TabForestPFN on the WhyTrees benchmark. Performance of different ICL-transformers is often similar. All the benchmark algorithms are greyed out. See Appendix A.7 for colored benchmark lines.
s
Figure 9: Individual dataset results on the WhyTrees Benchmark. TabForestPFN shows the mean of ten default runs, all others use random search over the hyperparameters. See Table 2 for other ICL-transformers.
@@ A.9 Synthetic Data with Lower Complexity
Figure 11: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 32, dataset size 1024, tree depth between 1 and 25, two features, and between 2 and 10 number of classes. See also Figures 3 and 12.
Figure 12: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 1024, dataset size 1024, tree depth between 1 and 9, two features, and between 2 and 10 number of classes. See also Figures 3 and 11.
A.10 Decision Boundaries for Scaled TabForest
In the complexity boundary results, we saw that TabForest has more complex decision boundaries when fine-tuned. Here we show that the complexity of the decision boundary is correlated with the dimension size of the transformer. See Figure 13 for the results. We see that the decision boundaries become more complex when the dimension of the ICL-transformer increases.
Figure 13: Comparison of decision boundaries for the electricity dataset. Axis represent features, color is predicted class probability, dots are test observations. When increasing the dimension size of the transformer, complexity increases.

### v2

@@ A.6 TabZilla Further Results
In the TabZilla main results Table 3, we have shown the performance including all methods implemented by the TabZilla authors. Because the rank is calculated over all included methods, which ICL-transformer variants we include might change the results. Therefore, we check if the results are the same if we use calculate the rankings one ICL-transformer at the time.
@@
TabForestPFN - Fine-tuned 1 19 5.6 4.5 0.846 0.910
CatBoost 1 15 6.2 5.0 0.848 0.876
XGBoost 1 16 6.3 5.0 0.841 0.901
LightGBM 1 19 7.7 6.0 0.792 0.871
RandomForest 1 18 7.9 8.0 0.797 0.852
NODE 1 19 8.4 8.0 0.754 0.839
Resnet 1 19 8.4 8.0 0.729 0.837
SAINT 1 19 8.5 8.0 0.733 0.817
SVM 1 18 8.6 8.0 0.713 0.801
FT-Transformer 1 16 8.9 8.5 0.737 0.805
DANet 2 19 10.0 10.0 0.721 0.768
MLP-rtdl 1 19 11.5 12.0 0.622 0.737
STG 1 19 11.5 12.0 0.594 0.672
LinearRegression 1 19 12.3 13.8 0.567 0.592
MLP 1 19 12.6 14.0 0.572 0.588
TabNet 2 19 12.8 13.2 0.583 0.672
DecisionTree 1 19 13.3 14.0 0.504 0.551
KNN 2 19 13.9 15.0 0.475 0.484
VIME 1 19 15.7 17.0 0.345 0.240
@@
Zero-shot
TabScratch 19 19 19.0 19.0 0.000 0.000
TabPFN (original) 1 19 7.7 7.0 0.780 0.841
TabPFN (retrained) 1 18 7.0 6.0 0.803 0.860
TabForest 1 19 9.8 10.0 0.714 0.822
TabForestPFN 1 18 6.3 6.0 0.824 0.900
Fine-tuned
TabScratch 1 18 8.3 7.0 0.757 0.819
TabPFN (original) 1 18 6.4 6.5 0.838 0.909
TabPFN (retrained) 1 18 5.6 5.0 0.847 0.891
TabForest 1 19 7.0 6.0 0.810 0.885
TabForestPFN 1 19 5.6 4.5 0.846 0.910
@@ A.7 WhyTrees Further Results
The main results in Figure 4 report the normalized accuracy aggregated over all datasets. In Figure 8 and 9 we show the comparison between fine-tuned TabForestPFN and the original zero-shot version of TabPFN on all 23 datasets. In Figure 10 and 11 we show the same graphs but with fine-tuned TabPFN and fine-tuned TabForest.
Figure 8: Comparison of fine-tuned TabForestPFN and the original zero-shot TabPFN on the WhyTrees benchmark with mixed features.
Figure 9: Comparison of fine-tuned TabForestPFN and the original zero-shot TabPFN on the WhyTrees benchmark with mixed features.
Figure 10: Comparison of fine-tuned TabForest and fine-tuned TabPFN on the WhyTrees benchmark with mixed features.
Figure 11: Comparison of fine-tuned TabForest and fine-tuned TabPFN on the WhyTrees benchmark with mixed features.
@@ A.9 Synthetic Data with Lower Complexity
Figure 13: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 32, dataset size 1024, tree depth between 1 and 25, two features, and between 2 and 10 number of classes. See also Figures 3 and 14.
Figure 14: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 1024, dataset size 1024, tree depth between 1 and 9, two features, and between 2 and 10 number of classes. See also Figures 3 and 13.